# LBMamba: Locally Bi-directional Mamba

### v1

@@ Abstract
Mamba, a State Space Model (SSM) that accelerates training by recasting recurrence as a parallel selective scan, has recently emerged as a linearly-scaling, efficient alternative to self-attention. Because of its unidirectional nature, each state in Mamba only has information of its previous states and is blind to states after. Current Mamba-based computer-vision methods typically overcome this limitation by augmenting Mamba’s global forward scan with a global backward scan, forming a bi-directional scan that restores a full receptive field. However, this operation doubles the computational load, eroding much of the efficiency advantage that originally Mamba have. To eliminate this extra scans, we introduce LBMamba, a locally bi-directional SSM block that embeds a lightweight locally backward scan inside the forward selective scan and executes it entirely in per-thread registers. Building on LBMamba, we present LBVim, a scalable vision backbone that alternates scan directions every two layers to recover a global receptive field without extra backward sweeps. We validate the versatility of our approach on both natural images and whole slide images (WSIs). We show that our LBVim constantly offers a superior performance–throughput trade-off. That is under the same throughput, LBVim achieves 0.8% to 1.6% higher top-1 accuracy on the ImageNet-1K classification dataset, 0.6% to 2.7 % higher mIoU on the ADE20K semantic segmentation dataset, 0.9% higher APb and 1.1% higher APm on the COCO detection dataset. We also integrate LBMamba into the SOTA pathology multiple instance learning (MIL) approach, MambaMIL, which uses single directional scan. Experiments on 3 public WSI classification datasets for show that our method achieves a relative improvement of up to 3.06% better AUC, 3.39% better F1, 1.67% better accuracy.
@@ 1 Introduction
State-space models (SSMs) have emerged as a compelling alternative to self-attention for sequence modeling because their hidden-state recurrence yields linear time and memory complexity with respect to sequence length (Gu et al., 2021a; Wang et al., 2022). Yet, conventional SSMs trained with naïve recurrence is still limited by slow training and inference, as they cannot leverage efficient parallelism of modern GPUs (Baron et al., 2023). Mamba (Gu & Dao, 2023) overcomes this by decoupling the state update from the hidden-to-output convolution and reformulating the computation as a selective parallel scan that runs efficiently on modern GPUs. Consequently, Mamba matches Transformer-level accuracy on long-range tasks while exhibiting far better resolution wise scaling characteristics, making it an attractive choice for both research and production systems. It was first introduced for natural-language processing and has since been adapted to computer vision (Zhu et al., 2024; Liu et al., 2024; Huang et al., 2024). Vision models built on Mamba’s selective-scan kernel delivers substantial GPU speed-ups and memory savings while consistently outperforming Transformer based baselines.
Standard computer vision mamba based models scan images multiple times from different directions to enhance their performance (Zhu et al., 2024; Liu et al., 2024; Yang et al., 2024a). There are two causes of such multiple scans: the first one is to overcome the 1D nature of Mamba. Mamba treats an image as a flattened 1D sequence, so a single left-to-right pass captures only row-wise context. To recover vertical dependencies, vision pipelines typically perform an additional scan on the column-wise ordering of patches, yielding two orthogonal sweeps that together approximate 2D spatial relationship (Liu et al., 2024; Yang et al., 2024a). Several dedicated 2D Mamba/SSM methods have recently been proposed to address this structural mismatch more directly (Zhang et al., 2024; Wang et al., 2024). The second issue is the unidirectional nature of SSMs: the latent state at position $t$ is conditioned only on past positions $1\dots t$ (Gu & Dao, 2023). Consequently, the model is blind to information occurring after position $t$, which often leads to sub-optimal performance on vision tasks. A common solution is to add a reverse (right-to-left or bottom-to-top) pass to restore access to future tokens, producing a bi-directional scan mechanism (Zhu et al., 2024). Although this strategy re-establishes a full receptive field, each extra sweep roughly doubles the computational load, eroding much of the efficiency advantage that originally Mamba offered.
@@
We show that instead of conducting an extra backward scan, it is more beneficial to scale up the model size under a fixed latency budget. Experiments on natural images show that LBVim achieves better performance-throughput trade-off than the baselines. Under the same throughput, LBVim achieves 0.8% to 1.6% higher top-1 accuracy on the ImageNet-1K classification dataset, 0.6% to 2.7 % higher mIoU on the ADE20K semantic segmentation dataset, 0.9% higher APb and 1.1% higher APm on the COCO detection dataset. We also integrate our scanning approach into the SOTA pathology multiple instance learning (MIL) approach, MambaMIL pyang2024mambamil. Extensive experiments on 3 public datasets for WSI classification and survival analysis datasets show that our method achieves a relative improvement of up to $3.06\%$ better AUC, $3.39\%$ better F1, $1.67\%$ better accuracy.
@@ 2 Related work
State Space Model (SSM). SSM (Kalman, 1960) is an effective sequence model that represents systems evolving over time by defining hidden states and their transitions, which makes it particularly useful for capturing dynamic temporal behavior in sequential data. Gu et al. (2021b) unified RNNs, temporal convolutions, and neural differential equations with a linear state-space layer and demonstrated the potential of SSM-based models with the HiPPO initialization. Wang et al. (2022) proposed Bi-directional Gated SSM which is able to match BERT (Devlin et al., 2019) pretraining accuracy without attention. S4 (Gu et al., 2021a) normalized the parameter matrices into a diagonal structure and offered an option to use bi-directional convolution kernel. 2D-SSM (Baron et al., 2023) adopted a 2D-SSM recursion (Kung et al., 1977) and explored scanning image in two or four directions. Similarly, S4ND (Nguyen et al., 2022) extends S4 to images and videos by applying axis-wise updates successively along each spatial direction. Overall, SSM architectures developed prior to Mamba were constrained by slow training efficiency because there lacks an efficient parallel algorithm.
@@
Mamba. Mamba (Gu & Dao, 2023) proposed a selective mechanism that makes the model parameters input-dependent and thus allows remembering important states and discard less relevant ones, alleviating the forgetting in long sequences. It also introduced a hardware-aware parallel scan algorithm that drastically accelerates state computation. Hwang et al. (2024) proposed Hydra, a bi-directional Mamba model that using generalized matrix mixers and showed a better performance than BERT. Vim (Zhu et al., 2024) and VMamba (Liu et al., 2024) are the first two mamba based models in computer vision. Vim (Zhu et al., 2024) introduced a Vision Mamba block that uses two independent selective SSMs for bi-directional aggregation of information and achieves a global receipt field. VMamba (Liu et al., 2024) introduced a pyramid Mamba network with a 4-directional scan pattern to achieve a global receipt field and also enhance spatial understanding. It also proposed some solution to alleviate the instability issue in half precision training. PlainMamba (Yang et al., 2024a) also used a 4-directional selective scan and adopted a more spatially continuous scan path. 2DMmaba (Zhang et al., 2024) introduced a 2D SSM and extends Mamba’s parallel scan algorithm into this 2D SSM. It also use the 4-directional scan of VMamba in their natural image applications. For all of these 4-directional scans, they are all two groups of bi-directional scans applied on two different ordering of image patches. This bi-directional scan is now the standard approach in most mamba based vision models.
@@
Application of Mamba in Whole Slide Images (WSI) analysis. WSIs are usually Giga-pixel images in the pathology domain. Most slides annotated only on slide level, requiring Multi Instance Learning (MIL) methods for WSI classification. It aggregates embedded features from a WSI for slide-level representation. MIL approaches are usually based on some attention mechanism (Ilse et al., 2018; Lu et al., 2021; Li et al., 2021) and self-attention mechanism (Shao et al., 2021). Recently, S4MIL (Fillioux et al., 2023) introduced S4 model to WSI analysis, which demonstrated the effectiveness of SSM in capturing long-range dependencies, but it does not utilize the parallel scan and thus slow. Yang et al. (Yang et al., 2024b) used Mamba and a sequence reordering mechanism to reduce overfitting and achieved even better performance. Note that these SSM based MIL methods does not use bi-directional scan.
@@ 3.1 SSM in Mamba and existing standard bi-directional SSM
The original SSM in Mamba (Gu & Dao, 2023) is a mathematical model used to capture the behavior of continuous dynamic systems. To be integrated into deep models, it is discretized as:
@@
where $x_{t}$ is the input token, $h_{t}^{f}$ is the latent state at time $t$, $y_{t}^{f}$ is the output token, $D$ is a parameter, $C^{f}$ is the state dimension coefficient to aggregate $N$ state dimensions into a single output and we use the superscript $f$ to denote this is a forward scan instead of a backward one. In this paper, following previous conversion (Zhu et al., 2024), "forward" and "backward" represents the scan direction rather than the forward/backward propagation in neural network training. The parameters $\bar{A}^{f}$, $\bar{B}^{f}$ and $\bar{C}^{f}$ are functions of the input $x_{t}$, which allows the SSM to dynamically adapt to the input context (known as the selective mechanism (Gu & Dao, 2023)). This aggregates important input into the hidden state while unimportant input can be ignored. Mathematically, they are:
@@
From equation 1, we can find that the output $y_{t}^{f}$ only contains information of its previous inputs $\{x_{i}|i<t\}$. It is also illustrated in figure 1 left. In order to enable a global recept field, previous approaches (Zhu et al., 2024; Liu et al., 2024; Yang et al., 2024a) commonly apply a backward scan and add them together as the output $y_{t}$ (figure 1 ceneter):
@@
Figure 1: (Left): The unidirectional scanning mechanism of the vanilla Mamba (Gu & Dao, 2023), where each state only has information of its previous/left states. (Center): Standard bi-directional scanning mechanism (Zhu et al., 2024), where a dedicated backward scan is conducted and added to the forward scan. This operation involves an additional read/write of the data and thus doubles the running time. (Right): Our LBMamba scan conducts a locally backward scan which is integrated into the forward scan process. This involves only one time read/write operations and thus very fast.
@@ 3.2 Locally bi-directional SSM architecture
As shown in figure 1 right, we first conduct a global forward scan as equation 1. Then we divide the input sequence of length $L$ into sub-sequences which all have a length of $M$ ($M$=3 in the figure). $M$ is set to the number of elements one thread processes and we detail it section 3.4. In each of these sub-sequences, we conduct a backward scan as that in equation 4. Specifically, the state $h_{t}^{b}$ obtained during the local backward scan is:
@@ 3.3 Architecture of LBVim
Building upon LBMamba, we introduce the overall architecture of LBVim. LBVim is based on Vim (Zhu et al., 2024). As shown in figure 2, The patch embedding and position embedding follow the same design as Vim. The embedded patches are then processed by $U$ LBMamba encoders. We replace Vim’s global bi-directional selective scan with our locally bidirectional variant. As LBMamba  is not a global bi-directional scan, in one encoder each token does not have a global view of the sequence. To alleviate this problem, we reverse the feature sequence at the end of each encoder to alternate the global scan direction (forward or backward). It allows each token to achieve a global receptive field after every two encoders.
Notably, we does not use the class token commonly used in (Dosovitskiy et al., 2020; Zhu et al., 2024), as it underperforms in our model. Instead, we employ an Global Average Pooling (GAP) layer followed by an MLP for final prediction in small sized models. When the size of the model scales up, we use an Multi-Head Attention with latent query mechanism (MHA) to aggregate features from the last LBMamba encoder:
$class=MLP(softmax(\frac{qK^{T}(ft)}{\sqrt{d}}V(ft)))$ (9)
where $q$ is a single learnable token, similar to the class token, $ft$ are the features from the last LBMamba encoder, $d$ is the dimension of features, $K(\cdot),V(\cdot)$ are learnable linear functions and $class$ is the final prediction. Unlike the self attention mechanism, this attention only has one query and thus has linear time complexity. We find that on larger models, MHA achieves better performance than GAP.
@@ 3.4 Hardware-aware thread-level bi-directional scanning operator
GPU storage hierarchy. Figure 3 right illustrates the storage hierarchy of modern GPUs. The green region denotes off-chip GPU memory, with low speed and high capacity. It is referred to as high bandwidth memory (HBM). The orange area denotes on-chip static RAM (SRAM), with high speed but low capacity. The blue region highlights the per-thread registers, the fastest tier yet restricted to at most 255 registers per thread. In typical GPU algorithms, data is transferred from HBM to registers for computation, and the results are stored back to HBM to free SRAM and registers for succeeding computation. Because registers are private to each thread, inter-thread communication usually traverse the SRAM, which is substantially more costly than intra-thread computation. Large-scale HBM operations are even more expensive. As a consequence, many GPU algorithms (Dao et al., 2022; Dao, 2024), including Mamba, are bounded not by arithmetic computation but by memory bandwidth.

### v2

@@ Abstract
Mamba, a State Space Model (SSM) that accelerates training by recasting recurrence as a parallel selective scan, has recently emerged as a linearly-scaling, efficient alternative to self-attention. Because of its unidirectional nature, each state in Mamba only has information of its previous states and is blind to states after. Current Mamba-based computer-vision methods typically overcome this limitation by augmenting Mamba’s global forward scan with a global backward scan, forming a bi-directional scan that restores a full receptive field. However, this operation doubles the computational load, eroding much of the efficiency advantage that originally Mamba have. To eliminate this extra scans, we introduce LBMamba, a locally bi-directional SSM block that embeds a lightweight locally backward scan inside the forward selective scan and executes it entirely in per-thread registers. Building on LBMamba, we present LBVim, a scalable vision backbone that alternates scan directions every two layers to recover a global receptive field without extra backward sweeps. We validate the versatility of our approach on both natural images and whole slide images (WSIs). We show that our LBVim constantly offers a superior performance–throughput trade-off. That is under the same throughput, LBVim achieves 0.8% to 1.6% higher top-1 accuracy on the ImageNet-1K classification dataset, 0.6% to 2.7 % higher mIoU on the ADE20K semantic segmentation dataset, 0.9% higher APb and 1.1% higher APm on the COCO detection dataset. Our method serves as a general-purpose enhancement, boosting the accuracy of four SOTA Mamba models, namely VMamba, LocalVim, PlainMamba and Adventurer, by 0.5% to 3.4%. We also integrate LBMamba into the SOTA pathology multiple instance learning (MIL) approach, MambaMIL, which uses single directional scan. Experiments on 3 public WSI classification datasets for show that our method achieves a relative improvement of up to 3.06% better AUC, 3.39% better F1, 1.67% better accuracy. Our code is available at https://github.com/cvlab-stonybrook/LBMamba.
@@ 1 Introduction
State-space models (SSMs) have emerged as a compelling alternative to self-attention for sequence modeling because their hidden-state recurrence yields linear time and memory complexity with respect to sequence length (gu2021efficiently; wang2022pretraining). Yet, conventional SSMs trained with naïve recurrence is still limited by slow training and inference, as they cannot leverage efficient parallelism of modern GPUs (baron20232_2dssm). Mamba (gu2023mamba) overcomes this by decoupling the state update from the hidden-to-output convolution and reformulating the computation as a selective parallel scan that runs efficiently on modern GPUs. Consequently, Mamba matches Transformer-level accuracy on long-range tasks while exhibiting far better resolution wise scaling characteristics, making it an attractive choice for both research and production systems. It was first introduced for natural-language processing and has since been adapted to computer vision (vim; liu2024vmamba; huang2024localmamba). Vision models built on Mamba’s selective-scan kernel delivers substantial GPU speed-ups and memory savings while consistently outperforming Transformer based baselines.
Standard computer vision mamba based models scan images multiple times from different directions to enhance their performance (vim; liu2024vmamba; yang2024plainmamba). There are two causes of such multiple scans: the first one is to overcome the 1D nature of Mamba. Mamba treats an image as a flattened 1D sequence, so a single left-to-right pass captures only row-wise context. To recover vertical dependencies, vision pipelines typically perform an additional scan on the column-wise ordering of patches, yielding two orthogonal sweeps that together approximate 2D spatial relationship (liu2024vmamba; yang2024plainmamba). Several dedicated 2D Mamba/SSM methods have recently been proposed to address this structural mismatch more directly (zhang20242dmamba; wang2024v2m). The second issue is the unidirectional nature of SSMs: the latent state at position $t$ is conditioned only on past positions $1\dots t$ (gu2023mamba). Consequently, the model is blind to information occurring after position $t$, which often leads to sub-optimal performance on vision tasks. A common solution is to add a reverse (right-to-left or bottom-to-top) pass to restore access to future tokens, producing a bi-directional scan mechanism (vim). Although this strategy re-establishes a full receptive field, each extra sweep roughly doubles the computational load, eroding much of the efficiency advantage that originally Mamba offered.
@@
We show that instead of conducting an extra backward scan, it is more beneficial to scale up the model size under a fixed latency budget. Experiments on natural images show that LBVim achieves better performance-throughput trade-off than the baselines. Under the same throughput, LBVim achieves 0.8% to 1.6% higher top-1 accuracy on the ImageNet-1K classification dataset, 0.6% to 2.7 % higher mIoU on the ADE20K semantic segmentation dataset, 0.9% higher APb and 1.1% higher APm on the COCO detection dataset. Applying our approach to four diverse state-of-the-art (SOTA) models: VMamba (liu2024vmamba), LocalVim (huang2024localmamba), PlainMamba (yang2024plainmamba), and Adventurer (wang2025adventurer), improves their accuracy by 0.5% to 3.4%, demonstrating its effectiveness as a general-purpose enhancement for Mamba-based models. We also integrate our scanning approach into the SOTA pathology multiple instance learning (MIL) approach, MambaMIL (yang2024mambamil). Extensive experiments on 3 public datasets for WSI classification and survival analysis datasets show that our method achieves a relative improvement of up to $3.06\%$ better AUC, $3.39\%$ better F1, $1.67\%$ better accuracy.
@@ 2 Related work
State Space Model (SSM). SSM (kalman1960new_ssm) is an effective sequence model that represents systems evolving over time by defining hidden states and their transitions, which makes it particularly useful for capturing dynamic temporal behavior in sequential data. gu2021combining unified RNNs, temporal convolutions, and neural differential equations with a linear state-space layer and demonstrated the potential of SSM-based models with the HiPPO initialization. wang2022pretraining proposed Bi-directional Gated SSM which is able to match BERT (devlin2019bert) pretraining accuracy without attention. S4 (gu2021efficiently) normalized the parameter matrices into a diagonal structure and offered an option to use bi-directional convolution kernel. 2D-SSM (baron20232_2dssm) adopted a 2D-SSM recursion (kung1977new_2dssm_theory) and explored scanning image in two or four directions. Similarly, S4ND (nguyen2022s4nd) extends S4 to images and videos by applying axis-wise updates successively along each spatial direction. Overall, SSM architectures developed prior to Mamba were constrained by slow training efficiency because there lacks an efficient parallel algorithm.
@@
Mamba. Mamba (gu2023mamba) proposed a selective mechanism that makes the model parameters input-dependent and thus allows remembering important states and discard less relevant ones, alleviating the forgetting in long sequences. It also introduced a hardware-aware parallel scan algorithm that drastically accelerates state computation. hwang2024hydra proposed Hydra, a bi-directional Mamba model that using generalized matrix mixers and showed a better performance than BERT. Vim (vim) and VMamba (liu2024vmamba) are the first two mamba based models in computer vision. Vim (vim) introduced a Vision Mamba block that uses two independent selective SSMs for bi-directional aggregation of information and achieves a global receipt field. VMamba (liu2024vmamba) introduced a pyramid Mamba network with a 4-directional scan pattern to achieve a global receipt field and also enhance spatial understanding. It also proposed some solution to alleviate the instability issue in half precision training. EfficientVMamba (pei2024efficientvmamba) introduced skip sampling in scanning and replace the majority of VMmaba layers to Inverted Residual blocks to speed up VMamba. Similarly, MSVMamba (shi2024multi_msvmamba) combines low resolution scales and high resolution scans to speed up VMamba. GrootVL xiao2024grootvl introduced a tree style scanning patten based on the minimum spanning tree, but it did not utilize the parallel scan of Mamba. Mamba-R (wang2025mamba_mamba_r) investigated adding regiters to the Mamba network. PlainMamba (yang2024plainmamba) used a 4-directional selective scan and adopted a more spatially continuous scan path. 2DMmaba (zhang20242dmamba) introduced a 2D SSM and extends Mamba’s parallel scan algorithm into this 2D SSM. It also use the 4-directional scan of VMamba in their natural image applications. For all of these 4-directional scans, they are all two groups of bi-directional scans applied on two different ordering of image patches. This bi-directional scan is now the standard approach in most mamba based vision models. A previous work similar to our approach is Adventurer (wang2025adventurer), where the sequence is also reversed every consecutive layers to remove the bi-directional scans. Adventurer enhances the global understanding by prepending an average-pooled token while our method focuses on the local understanding of the model.
@@
Application of Mamba in Whole Slide Images (WSI) analysis. WSIs are usually Giga-pixel images in the pathology domain. Most slides annotated only on slide level, requiring Multi Instance Learning (MIL) methods for WSI classification. It aggregates embedded features from a WSI for slide-level representation. MIL approaches are usually based on some attention mechanism (ilse2018attention_abmil; lu2021data_clam; li2021dual_dsmil) and self-attention mechanism (shao2021transmil). Recently, S4MIL (fillioux2023structured_s4mil) introduced S4 model to WSI analysis, which demonstrated the effectiveness of SSM in capturing long-range dependencies, but it does not utilize the parallel scan and thus slow. Yang et al. (yang2024mambamil) used Mamba and a sequence reordering mechanism to reduce overfitting and achieved even better performance. M3Mamba (zheng2025m3amba) used dynamic memory bank to overcome the catastrophic forgetting of Mamba for long-term context. Note that these SSM based MIL methods does not use bi-directional scan.
@@ 3.1 SSM in Mamba and existing standard bi-directional SSM
The original SSM in Mamba (gu2023mamba) is a mathematical model used to capture the behavior of continuous dynamic systems. To be integrated into deep models, it is discretized as:
@@
where $x_{t}$ is the input token, $h_{t}^{f}$ is the latent state at time $t$, $y_{t}^{f}$ is the output token, $D$ is a parameter, $C^{f}$ is the state dimension coefficient to aggregate $N$ state dimensions into a single output and we use the superscript $f$ to denote this is a forward scan instead of a backward one. In this paper, following previous conversion (vim), "forward" and "backward" represents the scan direction rather than the forward/backward propagation in neural network training. The parameters $\bar{A}^{f}$, $\bar{B}^{f}$ and $\bar{C}^{f}$ are functions of the input $x_{t}$, which allows the SSM to dynamically adapt to the input context (known as the selective mechanism (gu2023mamba)). This aggregates important input into the hidden state while unimportant input can be ignored. Mathematically, they are:
@@
From equation 1, we can find that the output $y_{t}^{f}$ only contains information of its previous inputs $\{x_{i}|i<t\}$. It is also illustrated in figure 1 left. In order to enable a global recept field, previous approaches (vim; liu2024vmamba; yang2024plainmamba) commonly apply a backward scan and add them together as the output $y_{t}$ (figure 1 ceneter):
@@
Figure 1: (Left): The unidirectional scanning mechanism of the vanilla Mamba (gu2023mamba), where each state only has information of its previous/left states. (Center): Standard bi-directional scanning mechanism (vim), where a dedicated backward scan is conducted and added to the forward scan. This operation involves an additional read/write of the data and thus doubles the running time. (Right): Our LBMamba scan conducts a locally backward scan which is integrated into the forward scan process. This involves only one time read/write operations and thus very fast.
@@ 3.2 Locally bi-directional SSM architecture
As shown in figure 1 right, we first conduct a global forward scan as equation 1. Then we divide the input sequence of length $L$ into sub-sequences which all have a length of $M$ (MM=3 in the figure). $M$ is set to the number of elements one thread processes and we detail it section 3.4. In each of these sub-sequences, we conduct a backward scan as that in equation 4. Specifically, the state $h_{t}^{b}$ obtained during the local backward scan is:
@@ 3.3 Architecture of LBVim
Building upon LBMamba, we introduce the overall architecture of LBVim. LBVim is based on Vim (vim). As shown in figure 2, The patch embedding and position embedding follow the same design as Vim. The embedded patches are then processed by $U$ LBMamba encoders. We replace Vim’s global bi-directional selective scan with our locally bidirectional variant. As LBMamba  is not a global bi-directional scan, in one encoder each token does not have a global view of the sequence. To alleviate this problem, we reverse the feature sequence at the end of each encoder to alternate the global scan direction (forward or backward). It allows each token to achieve a global receptive field after every two encoders.
Notably, we does not use the class token commonly used in (dosovitskiy2020image_vit; vim), as it underperforms in our model. Instead, we employ an Global Average Pooling (GAP) layer followed by an MLP for final prediction in small sized models. When the size of the model scales up, we use an Multi-Head Attention with latent query mechanism (MHA) to aggregate features from the last LBMamba encoder:
$class=MLP(softmax(\frac{qK^{T}(h^{U})}{\sqrt{d}}V(h^{U})))$ (9)
where $q$ is a single learnable token, similar to the class token, $h^{U}$ are the output tokens from the last LBMamba encoder, $d$ is the dimension of features, $K(\cdot),V(\cdot)$ are learnable linear functions and classclass is the final prediction. Unlike the self attention mechanism, this attention only has one query and thus has linear time complexity. We find that on larger models, MHA achieves better performance than GAP.
@@ 3.4 Hardware-aware thread-level bi-directional scanning operator
GPU storage hierarchy. Figure 3 right illustrates the storage hierarchy of modern GPUs. The green region denotes off-chip GPU memory, with low speed and high capacity. It is referred to as high bandwidth memory (HBM). The orange area denotes on-chip static RAM (SRAM), with high speed but low capacity. The blue region highlights the per-thread registers, the fastest tier yet restricted to at most 255 registers per thread. In typical GPU algorithms, data is transferred from HBM to registers for computation, and the results are stored back to HBM to free SRAM and registers for succeeding computation. Because registers are private to each thread, inter-thread communication usually traverse the SRAM, which is substantially more costly than intra-thread computation. Large-scale HBM operations are even more expensive. As a consequence, many GPU algorithms (dao2022flashattention; dao2023flashattention2), including Mamba, are bounded not by arithmetic computation but by memory bandwidth.